OpenStudy (anonymous):

I will award you! :) Evaluate the function for x=m/4 f(x)=8x^2-4x A. 2m^2-m B. m^2-m C. m^2/2-4m D. m^2/2-m

OpenStudy (anonymous):

@SolomonZelman :)

OpenStudy (anonymous):

= 8 * (m/4)^2 - 4(m/4) Simplify : 8 * (m^2/16) - m = m^2/2 - m :)

OpenStudy (anonymous):

hmmm. Thanks! :)